OpenStudy (anonymous):

solve 3/4 x -7 = 8 11.25 20 4/3 3/4

OpenStudy (anonymous):

@Kitty449

OpenStudy (anonymous):

Okay, instead of just giving you the answer, I can help you understand how to do it if you're willing to take the five or so minutes it will take.

OpenStudy (anonymous):

20

OpenStudy (anonymous):

its better if you tun the fraction into a decimal first. so it 0.75x-7=8 then you do the same thing we did in the last one.

OpenStudy (anonymous):

ok thanks. :D

OpenStudy (whpalmer4):

@kitty449 strongly disagree that you should turn fractions into decimals! What if you have a repeating decimal? What if I gave you a problem with \(1/97\) as a fraction? Even if you are going to write it with repeating decimal notation, there are 96 digits in the repeating portion! Also, converting to a decimal makes it more difficult to spot possible cancellations and factors. Do you recognize that 0.00952381... can be factored as \[\frac{1}{3}*\frac{1}{5}*\frac{1}{7}\]I'm guessing probably not. But it isn't hard to work that out from \(\dfrac{1}{105}\) which is the fraction whose decimal expansion starts with that string of digits. Work the problem without decimals: \[\frac{3}{4} x -7 = 8\] Add 7 to both sides: \[\frac{3}{4}x-7+7 = 8+7\]\[\frac{3}{4}x = 15\]Multiply both sides by 4\[4*\frac{3}{4}x = 4*15\]\[3x = 60\]divide both sides by 3\[\frac{3x}{3} = \frac{60}3\]\[x=20\] Checking the answer: \[\frac{3}{4}x-7=8\]substitute our answer\[\frac{3}{4}(20)-7=8\]\[15-7=8\]\[8=8\checkmark\] True statement obtained by putting our answer in original equation, so our answer is correct. No muss, no fuss, no need for any calculating devices.